OpenStudy (anonymous):

Can somebody help me to explain, in complete sentences, how you would I use the elimination method to solve the following system of equations. and also give me the solution? thank you 6x – 5y = 4 9x + 4y = –17

OpenStudy (anonymous):

eliminate y from second equation

OpenStudy (anonymous):

can you please explain that to me so I can write it? Thanks

OpenStudy (anonymous):

this is what I wrote :"I can use the elimination method to solve system equations because if I know how to solve for one variable i can get rid of the other one. " But I still dont know the asnwer to the problem..

OpenStudy (lilg132):

first you would need to make one of the variables the same value we will use the y variable for example only because it would be easier as one of them is already a negative and one is a positive so to make them both the same value you must multiply the two equations by something 6x – 5y = 4 (multiply by 4) 9x + 4y = –17 (multiply by 5) which will give you 24x - 20y = 16 45x + 20y = -85 add both the equations together the y variables will cancel eachother out because they are now the same value and when you add them together they will cancel eachother out because one is a negative and one is a positvive so you will now get 69x = -69 now to get the x variable on its own you must divide both sides by the number next to x 69x / 69 = -69 / 69 x = -1 now we know what x is we can use this to work out what the y variable is so we go back to our original equations and choose either one for this example i will choose the first equation 6x – 5y = 4 now everywhere you see the variable x you replace it with the value we have found which is -1 so the equation will now look like this 6(-1) -5y = 4 which is -6 -5y = 4 rearrange this so we have the variable on one side and the numbers on the other remember when bringin a variable or a number to the other side will change its value to the opposite e.g. bringin -6 to the other side will make it +6 so you now have -5y = 4 + 6 -5y = 10 -5y / 5 = 10 / 5 -y = 2 now to change the y into positive we swap the sides around turning the y into a positive and the number 2 into a negative -2 = y put the two figures we have found in the formula to check if you get the same answer. remember to click good answer

OpenStudy (anonymous):

thank you.. and I did